2^(x+1) * 3^(y+1)=(2*3)^5 Find 2x+y Please give the solution

  • 2^(x+1) * 3^(y+1)=(2*3)^5
  • Find 2x+y
  • Please give the solution

Grade:11

2 Answers

Rinkoo Gupta
9 years ago
comparing theindices of 2 and 3 on boths sides
x+1=5, y+1=5
x=4, y=4
hence 2x+y=2(4)+4=12
Thanks & Regards
Rinkoo Gupta
AskIITians Faculty
Sher Mohammad IIT Delhi
9 years ago
2^(x+1)*3^(y+1)=2^5*3^5
2 and 3 are coprime so no common factor between them so
x+1=5=>x=4
y+1=5=.y=4
so 2x+y=12

Sher Mohammad
faculty askiitians